Find the price elasticity of demand at the point where quantity demanded is equal to quantity supplied


1
Expert's answer
2021-03-12T07:25:10-0500

The price elasticity of demand at the equilibrium can be found as follows:


"E_d=(-\dfrac{1}{slope})\cdot\dfrac{P}{Q_d},"


here, "P" is the equilibrium price and "Q_d" is the equilibrium quantity.